There were 587 tickets purchased for a major league baseball game. The lower box  tickets cost ​$12.50 and the upper reserved tickets cost ​$8.00. The total amount of money spent was ​$6046.00. How many of each kind of ticket were​ purchased?
Answer by addingup(3677)

lower : x
upper : y
x + y = 587 subtract y from both sides and you get x = 587 - y We'll use this value for x next.
12.50x + 8.00y = 6,046.00
Now substitute for x:
12.50(587 - y) + 8.00y = 6,046.00
7,337.50 - 12.50y + 8.00y = 6,046.00
-4.50y = -1291.50 divide to find y, and remember -/- = +
y = 287
287 of the $8 tickets were purchased, and 587-287 = 300 of the 12.50 tickets were purchased